# Near–Real-Time Conflict-Related Fire Detection Using Unsupervised Deep Learning and Satellite Imagery

### v2

@@ 5.2 Observations on False Positives and False Negatives
Analysis of tile-level outputs revealed patterns in prediction errors. False positives were most commonly associated with residual burned material from prior incidents or structural debris resembling charred surfaces, reflecting the model’s sensitivity to broader damage signatures. False negatives occurred primarily for small, fragmented fires below the 3 m sensor resolution in densely built urban areas. These error modes are consistent with known limitations of optical imagery in conflict environments and are considered acceptable within operational settings that prioritize rapid situational awareness over exhaustive error minimization (Zwijnenburg and Ballinger, 2023). Despite these challenges, recall remained relatively high, consistent with the near–real-time monitoring objective that prioritizes high-sensitivity detection of fire-affected areas.
@@ 5.3 Sufficiency of 4-band Imagery
Across all locations, 4-band PlanetScope imagery captured most discriminative features for fire detection. While modest improvements were observed with 8-band imagery or short temporal sequences (median relative AUPRC gains of 3–7% across tested scenes), statistical testing indicated that these improvements were small and often not operationally significant. We note that this observation is specific to the scenes and fire types analyzed; additional spectral bands may improve performance in areas with more subtle spectral signatures of fire or more heterogeneous land cover. In El Fasher, the modest improvement observed with 8-band imagery, and the degraded baseline performance, are likely due to increased sensitivity to inter-band noise and residual misregistration in a highly heterogeneous urban setting.
@@ 5.4 Summary
Overall, the proposed VAE-based approach demonstrates robust and statistically significant improvements over a pixel-space baseline. Bootstrapped confidence intervals confirm stability across scenes, effect sizes indicate substantial performance gains, and observed error patterns provide actionable insight for operational near–real-time monitoring of conflict-related fires.
All experiments are compatible with the near–real-time processing pipeline described in Section 4.3, producing outputs within 24–30 hours of image acquisition. Consistent with prior work (Růžička et al., 2022), these results demonstrate that latent representation–based change detection substantially outperforms direct pixel-space comparisons. Considering the acquisition, preprocessing, and computational overheads of multi-spectral and multi-temporal inputs, the limited performance gains observed with additional data modalities indicate that the proposed lightweight 4-band configuration offers a favorable trade-off between accuracy and operational efficiency for near–real-time conflict monitoring. We note that all quantitative results are conditioned on the availability of cloud-free optical imagery and accurate co-registration between pre- and post-incident scenes.
@@ 6 Discussion
Our findings demonstrate that near–real-time detection of conflict-related fire damage is feasible using near-daily PlanetScope imagery combined with latent representation–based change detection. Across multiple case studies, the proposed VAE-based approach outperformed a pixel-space cosine distance baseline, producing spatially coherent outputs suitable for rapid situational awareness. These findings align with prior research demonstrating the potential of latent representation learning in satellite-based change detection (Bennett et al., 2022). Additionally, results indicate that 4-band imagery captures most of the discriminative information required for fire detection, with additional spectral bands or short temporal sequences providing only modest gains. Compared with public satellites such as MODIS, VIIRS, and Sentinel-2, PlanetScope provides a practical balance of spatial resolution, revisit frequency, and accessibility. However, its limited 4–8 band coverage constrains detection of subtle spectral signatures compared with Sentinel-2’s 13 bands. This trade-off reflects methodological choices common in conflict monitoring: 3 m resolution commercial imagery allows near-daily monitoring at manageable cost, while very high-resolution imagery remains limited due to availability and expense (Bennett et al., 2022). This work highlights the importance of interdisciplinary integration. Verification of detected incidents, uncertainty quantification, and operational relevance require input from peace and conflict research, humanitarian studies, and remote sensing expertise. Prior work emphasizes that automated satellite detection alone does not produce actionable insight; coordinated approaches are needed to interpret results, ensure reliability, and support evidence-based decision-making in conflict-affected contexts (Ghioni et al., 2024; Duursma, 2023).
@@ 7 Conclusion
A civil war in Sudan, primarily between the Sudan Armed Forces (SAF) and the Rapid Support Forces (RSF), has caused widespread civilian casualties and displacement (Ahmed et al., 2025), with over 1,400 violent incidents targeting civilians, resulting in 61,000 war-related deaths and 12 million displaced people (Rothbart et al., 2025). This study demonstrates that conflict-related fire damage in Sudan can be detected in near–real-time using PlanetScope imagery and latent representation–based change detection. The VAE-based approach reliably detected fire damage, with 4-band imagery providing the most relevant information and additional bands or short temporal sequences offering only minor improvements. The study contributes an operationally focused integration of remote sensing, machine learning, and conflict analysis to support humanitarian assessment in data-scarce regions. We note some limitations of the work: it is based on a small number of case studies in Sudan, which limits generalizability, the baseline comparison relies on a simple cosine distance method, and dependence on commercial PlanetScope imagery introduces constraints on data access, cost, and reproducibility. Despite these limitations, the results demonstrate the potential of lightweight, latent representation–based methods for near–real-time monitoring of conflict-related fire damage, illustrating a practical trade-off between accuracy, data availability, and operational efficiency for rapid conflict monitoring when higher-resolution or ground-based data are unavailable.
@@ Acknowledgements
The code is available in a GitHub repository at https://github.com/heykuldip/nrt_conflict_monitoring

### v3

@@ 5.2 Observations on False Positives and False Negatives
Analysis of tile-level outputs revealed systematic patterns in prediction errors. False positives were primarily linked to residual burned material from prior incidents or debris resembling charred surfaces, reflecting the model’s sensitivity to broader damage signatures. Instances where active fires or burn scars went undetected were primarily restricted to small, fragmented events physically smaller than the 3 m spatial resolution of the PlanetScope sensor. These omissions are an inherent hardware limitation of optical remote sensing in densely built urban areas. These observational constraints align with known limitations of optical imagery in conflict environments, and are considered an acceptable trade-off in operational settings that prioritize rapid, broad-scale situational awareness over exhaustive ground-truth error minimization [67]. Despite these sensor-level challenges, overall model recall remained high, supporting the framework’s near–real-time monitoring objective of sensitive detection of fire-affected areas.
@@ 5.3 Sufficiency of 4-band Imagery
Across all locations, 4-band PlanetScope imagery captured sufficient spectral information for fire detection. While modest improvements were observed with 8-band imagery or short temporal sequences (median relative AUPRC gains of 3–7% across tested scenes), paired statistical tests indicated that these gains were small and often not operationally significant. We note that this observation is specific to the scenes and fire types analyzed; additional spectral bands may provide greater benefits in regions with more subtle fire signatures or heterogeneous land cover. In El Fasher, both the 8-band and time-series configurations yielded slightly lower AUPRC and Recall than the 4-band imagery, likely due to increased sensitivity to inter-band noise and residual misregistration errors in highly heterogeneous urban settings.
@@ 5.4 Summary
Overall, the proposed VAE-based approach indicates consistent and statistically significant improvements over CVA, cosine distance, and IR-MAD. Bootstrapped confidence intervals support stability across scenes, effect sizes suggest meaningful performance gains, and observed error patterns provide actionable insight for operational near–real-time monitoring of conflict-related fires. All experiments are compatible with the near–real-time processing pipeline described in Section 4.3, producing outputs within 24–30 hours of image acquisition. Consistent with prior work [56, 22], these results indicate that latent representation–based change detection tends to improve performance over conventional pixel-difference and transformation-based methods. Considering the acquisition, preprocessing, and computational overheads of multi-spectral and multi-temporal inputs, the limited performance gains observed with additional data modalities suggest that the proposed lightweight 4-band configuration offers a favorable trade-off between accuracy and operational efficiency for near–real-time conflict monitoring. We note that all quantitative results are conditioned on the availability of cloud-free optical imagery and accurate co-registration between pre- and post-incident scenes.
@@ 6 Discussion
Our findings demonstrate that near–real-time detection of conflict-related fire damage is feasible using near-daily PlanetScope imagery combined with latent representation–based change detection. Across multiple case studies, the proposed VAE-based approach yields improved performance relative to CVA, cosine distance, and IR-MAD, producing spatially coherent outputs suitable for rapid situational awareness. These findings align with prior research demonstrating the potential of latent representation learning in satellite-based change detection [9]. Additionally, results indicate that 4-band imagery captures sufficient discriminative information required for fire detection, with additional spectral bands or short temporal sequences providing only modest gains.
Furthermore, training on the ecologically distinct WorldFloods dataset reduces the risk of domain-specific overfitting, as the unsupervised VAE has no prior exposure to Sudan’s fire patterns or urban morphology. The system is intended to verify reported incidents, so evaluation on geolocated reports aligns with operational use, where focusing on known locations is an acceptable trade-off for timely validation in data-scarce regions.
Compared with public satellite platforms such as MODIS, VIIRS, and Sentinel-2, PlanetScope provides a practical balance between spatial resolution and revisit frequency that is particularly advantageous in rapidly evolving conflict environments. While Sentinel-2 offers 13 spectral bands, including SWIR channels commonly used for burn detection, its effective revisit interval (often 5–10 days) can limit timely assessment of rapidly changing fire events. PlanetScope’s near-daily 3 m imagery reduces temporal gaps and facilitates more continuous monitoring of localized damage patterns. However, this temporal advantage is accompanied by spectral limitations, as the absence of SWIR bands restricts the direct application of conventional burn indices and threshold-based approaches that rely on SWIR sensitivity to char and moisture loss [29].
This spatial–temporal–spectral trade-off motivates the adoption of an unsupervised variational autoencoder (VAE) framework. Rather than relying on predefined spectral indices or labeled training data, the VAE learns latent representations from the available RGB and Near-Infrared (NIR) bands, enabling detection of fire-related anomalies through changes in feature distributions between pre- and post-event imagery. In settings where timely updates are required and labeled datasets are limited, such an unsupervised approach leverages PlanetScope’s high revisit frequency while minimizing its spectral constraints. This methodological choice is therefore consistent with the operational objective of near–real-time conflict monitoring. By independently evaluating spatial partitions, the VAE framework has a linear time complexity with respect to the number of input tiles, making daily monitoring of massive regions like Darfur computationally feasible [45, 52].
This work highlights the importance of interdisciplinary integration. Verification of detected incidents, uncertainty quantification, and operational relevance require input from peace and conflict research, humanitarian studies, and remote sensing expertise. Prior work emphasizes that automated satellite detection alone does not produce actionable insight; coordinated approaches are needed to interpret results, ensure reliability, and support evidence-based decision-making in conflict-affected contexts [28, 23].
@@ 7 Conclusion
A civil war in Sudan, primarily between the Sudan Armed Forces (SAF) and the Rapid Support Forces (RSF), has caused widespread civilian casualties and displacement [2], with over 1,400 violent incidents targeting civilians, resulting in 61,000 war-related deaths and 12 million displaced people [55]. This study demonstrates that conflict-related fire damage in Sudan can be detected in near–real-time using PlanetScope imagery and latent representation–based change detection. The VAE-based approach demonstrated improved detection capabilities over traditional baselines, with 4-band imagery providing the most relevant information and additional bands or short temporal sequences yielding modest improvements.
This study presents an approach integrating remote sensing, unsupervised learning, and conflict analysis to support humanitarian assessment in data-scarce environments. We acknowledge several limitations of the work. The analysis is based on a limited number of case studies within Sudan, which may constrain generalizability. The comparative evaluation is restricted to established pixel-level change detection approaches grounded in spectral distance and transformation-based techniques, and reliance on commercial PlanetScope imagery introduces considerations related to data access, cost, and reproducibility.
Despite these limitations, the findings indicate that lightweight, latent representation–based methods offer a practical approach for near–real-time monitoring of conflict-related fire damage, highlighting a trade-off between spectral richness, temporal availability, and operational feasibility in settings where higher-resolution or ground-based data are scarce and operationally impractical to acquire in real-time.
@@ Acknowledgements
The code is available in a GitHub repository at https://github.com/heykuldip/nrt-conflict-monitoring